Rita Lamar and Jose sent a total of 54 text messages over their cell phone during the weekend. Rita sent 10 more than Lamar . Jose sent 2 times as many messages as Rita . How many did they send each
Answer by ikleyn(52909)

.
Rita Lamar and Jose sent a total of 54 text messages over their cell phone during the weekend.
Rita sent 10 more than Lamar . Jose sent 2 times as many messages as Rita .
How many did they send each
~~~~~~~~~~~~~~~~~~


    Rita   + Lamar + Jose     = 54

    (x+10) +   x   + 2*(x+10) = 54


Simplify and solve


    4x + 30 = 54

    4x      = 54 - 30 = 24

     x                = 24/4 = 6.


ANSWER.  Lamar sent 6;  Rita sent x+10 = 16;  Jose sent 2*16 = 32.


CHECK.  6 + 16 + 32 = 54.      ! Correct !

Solved.